# Fast nonparametric inference of network backbones for graph sparsification

### v1

@@ II.1 Global Network Backboning Algorithm
Our problem considers an input network $G=\{(i,j,w_{ij})\}$, which we represent as a set of $E$ tuples $(i,j,w_{ij})$ that indicate an edge from node $i$ to node $j$ with weight $w_{ij}$. $G$ may have self-edges or multi-edges—multi-edges can be combined into a single edge weight—but the weights $w_{ij}$ must be positive and integer-valued. $G$ may be directed or undirected, and in the undirected case we simply consider both $(i,j,w_{ij})$ and $(j,i,w_{ij})$ to be in $G$, with the edge $(i,j,w_{ij})$ appearing in an undirected backbone if either $(i,j,w_{ij})$ or $(j,i,w_{ij})$ is retained in the directed backbone (as done in serrano2009extracting ). Bipartite networks can also be accommodated by our framework. We will let $W=\sum_{(i,j,w_{ij})\in G}w_{ij}$ denote the total edge weight in the weighted graph $G$. The goal of our network backboning method is to extract a subset $G^{(b)}\subseteq G$ of weighted edges that best compress the overall weight distribution in $G$. We call the subset $G^{(b)}$ the backbone of $G$, which has $E^{(b)}\leq E$ edges with a total weight of $W^{(b)}\leq W$.
To find the optimal backbone $G^{(b)}$ corresponding to an input network $G$, we consider transmitting the weight structure in $G$ to a receiver using $G^{(b)}$ as an intermediate step. We assume the edge topology of $G$ is known by the receiver and we must transmit the edge weight $w_{ij}$ corresponding to every edge $(i,j)$. This is the same assumption made implicitly by many network backboning methods serrano2009extracting ; marcaccioli2019polya ; dianati2016unwinding , which consider the statistical significance of edge weights while conditioning on the degrees of nodes.
The first transmission scheme we consider encodes the backbone $G^{(b)}$ and the remaining non-backbone edges $G\setminus G^{(b)}$ in a global manner, accounting for the distribution of weights across all edges when choosing to keep an edge in the backbone $G^{(b)}$. In this encoding we must first specify the total weight $W$ of the graph. Since the receiver knows the topology of $G$, they know the total number of edges $E$, and since the network must have positive integer-valued weights they also know that each edge has a weight of at least $1$. To specify the total weight $W$ we can thus specify the excess weight $W-E$ beyond the minimum weight of $E$ if all edges had weight $w_{ij}=1$. To specify $W-E$, we must send a bitstring of length approximately $\log(W-E+1)$ bits (the binary representation of the integer $W-E$), where $\log\equiv\log_{2}$. We then must specify the total weight $W^{(b)}$ of the backbone and the total number of edges $E^{(b)}$ in the backbone, which must fall in the ranges $[0,W]$ and $[0,E]$ respectively so require $\log(W+1)$ and $\log(E+1)$ bits respectively to specify.
The next step is to transmit which edges $(i,j)$ are in the backbone. There are $E$ possible edges to choose from, and the receiver now knows there are $E^{(b)}$ edges in the backbone, meaning there are (${E\choose E^{(b)}}$ possibilities for the backbone edges. Therefore to specify the particular configuration among these possibilities we need a bitstring of length approximately $\log{E\choose E^{(b)}}$. Since the weight of each edge $(i,j)$ in the backbone will satisfy $w_{ij}\geq 1$, the list of weights for these edges will take the form of a length-$E^{(b)}$ positive integer vector that sums to $W^{(b)}$. The number of such vectors is equal to (${W^{(b)}-1\choose E^{(b)}-1}$, which can be computed using the standard “stars and bars” argument from combinatorics feller1991introduction . Therefore specifying the backbone weights requires $\log{W^{(b)}-1\choose E^{(b)}-1}$ additional bits. Finally, we specify the non-backbone edges $G\setminus G^{(b)}$ using the same approach, which through an analogous argument requires $\log{W-W^{(b)}-1\choose E-E^{(b)}-1}$ bits.
Figure 1: Global and local MDL backbones for a small example network. (a) Schematic of the global backbone transmission objective of Eq. II.1, with the total weight and number of edges indicated for the backbone $G^{(b)}$ (left) and remaining edges (right). (b) Schematic of the local backbone transmission objective of Eq. 5, applied to out-neighborhoods $\partial_{i}$ for each node $i$. The total strength and degree are indicated for the backbone neighborhood $\partial^{(b)}_{i}$ around an example node $i$. Each method aims to transmit a backbone that parsimoniously captures the weight structure in the network $G$ according to an information encoding and learns an optimal number of backbone edges nonparametrically. The local method provides a more compressive backbone for this network, giving an inverse compression ratio (Eq. 6) of $\eta^{(\text{local})}=0.86$ versus $\eta^{(\text{global})}=0.88$ for the global backbone. The backboning methods are applicable to any network with integer-valued edge weights, including directed or undirected networks and unipartite or bipartite networks. The global and local backboning methods are also generalizable to a broader set of weight distributions on the edges using an equivalent Bayesian model, as discussed in Sec. II.3.
Putting this all together, we have that the description length of the network weights in $G$ using a backbone $G^{(b)}$ is given by
$\mathcal{L}^{(\text{global})}(G^{(b)})$ $=\log(W-E+1)+\log(W+1)$
$+\log(E+1)+\log{E\choose E^{(b)}}$ (1)
$+\log{W^{(b)}-1\choose E^{(b)}-1}+\log{W-W^{(b)}-1\choose E-E^{(b)}-1}.$
This encoding scheme is considered “global” because the description length couples all edges together under the total backbone weight $W^{(b)}$ and the backbone edge count $E^{(b)}$. By minimizing $\mathcal{L}^{(\text{global})}(G^{(b)})$ over all possible backbones $G^{(b)}$, we obtain the optimal backbone $G_{\text{global}}^{(b)}$ according to this global encoding scheme.
For a fixed number of edges $E^{(b)}$ in the backbone $G^{(b)}$, the description length in Eq. II.1 is monotonically decreasing in $W^{(b)}$ for $W,E\gg 1$. We can show this by computing
$\mathcal{L}^{(\text{global})}(W^{(b)}+1,E^{(b)})-\mathcal{L}^{(\text{global})}(W^{(b)},E^{(b)})$
$=\log{\frac{W_{b}}{W_{b}-E_{b}+1}\frac{(W-W_{b})-(E-E_{b})}{W-W_{b}-1}},$ (2)
which is less than or equal to zero for
$\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}+\frac{2W^{(b)}-W-E^{(b)}+1}{WE}.$ (3)
For $W,E\gg 1$, the second term on the RHS vanishes and we have that for $\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}$, the description length in Eq. II.1 is monotonically decreasing in $W^{(b)}$ for fixed $E^{(b)}$. Therefore, under this condition, we have that at any value of $E^{(b)}$ the best edge to add to the backbone $G^{(b)}$ is the remaining edge $(i,j)$ with the highest weight $w_{ij}$.
The above argument implies that a greedy edge addition procedure for constructing the backbone will be optimal for $W,E\gg 1$, since the condition $\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}$ must always hold under such a procedure—this condition is equivalent to the condition $\frac{W^{(b)}}{E^{(b)}}\geq\frac{W}{E}$, or that the average weight of the backbone is at least the average weight of the network. We can therefore approximately minimize the description length objective in Eq. II.1 in the regime $W,E\gg 1$ with the following simple algorithm:
  1. 1.
    Initialize an empty backbone $G^{(b)}=\{\}$ with $E^{(b)}=W^{(b)}=0$ and sort the edges $(i,j,w_{ij})\in G$ in decreasing order of weight $w_{ij}$. Initialize an array DL to store the description length value for each value of $E^{(b)}$. Compute $\text{DL}[E^{(b)}=0]$ using Eq. II.1.
  2. 2.
    Iterate through the ordered edges in $G$, and for each edge $(i,j,w_{ij})$: increment $E^{(b)}\mathrel{+}=1$ and $W^{(b)}\mathrel{+}=w_{ij}$ and compute $\text{DL}[E^{(b)}]$ using Eq. II.1.
  3. 3.
    After adding the last edge from $G$, identify the value $E_{\text{global}}^{(b)}=\text{argmin}[\text{DL}]$ for which the description length was minimized, and set the MDL-optimal backbone $G^{(b)}_{\text{global}}$ to be the first $E_{\text{global}}^{(b)}$ edges in the ordered $G$.
The bottleneck of this algorithm is simply the initial step of ordering the edges in $G$ by weight, so the overall runtime complexity is $O(E\log E)$—only marginally slower than constructing the edge set $G$ itself.
We can see that this optimization procedure will construct backbones $G^{(b)}$ that concentrate as much weight $W^{(b)}$ as possible on $E^{(b)}$ edges, up to some point at which it is no longer information theoretically preferable to continue adding edges to the backbone. The final backbone $G^{(b)}_{\text{global}}$ is the most effective subset of network edges for compressing the weight distribution across all edges in the network under the specified encoding scheme.
It is evident from its formulation that the global MDL method described above is simply a principled, nonparametric way to identify a global threshold on the weights for constructing the backbone. Therefore this method is naturally limited for constructing network backbones when node strengths are heterogeneous—this is a major motivation for local methods such as the Disparity or Polya Filter serrano2009extracting ; marcaccioli2019polya . In such cases, the backbone $G^{(b)}$ is likely to be highly fragmented and have many degree-zero nodes as a result of weakly interacting nodes losing all of their adjacent edges. However, when node strengths are highly homogeneous, the global MDL backboning method may still be statistically justified from the perspective of data compression (we will see this in Sec. III). Therefore, for any backboning task one should use both the global method of this section and the local method presented in the next section, then compare their minimum description length values to choose the optimal MDL backbone for the dataset $G$.
II.2 Local Network Backboning Algorithm
To construct a local network backboning objective, we can apply a similar encoding scheme as above but at the level of individual node neighborhoods, which corrects for heterogeneity in the overall strengths of the nodes in a similar manner as other local backboning methods serrano2009extracting ; marcaccioli2019polya . For simplicity of presentation, we will define each node neighborhood $\partial_{i}$ as the set edges with node $i$ as the source node. But the method can be applied in an identical manner by considering all edges in which $i$ is the target node (i.e. considering in-neighborhoods instead of out-neighborhoods). For undirected networks, this distinction will make no difference since the edge directions are duplicated, so we can just use the out-neighborhoods. Since the method described has a simple closed form description length, one can choose the backbone to be constructed based on in- or out-neighborhoods based on which provides better data compression according to the final description length value.
Let $\partial_{i}=\{(i,j):w_{ij}>0\}$ be the neighborhood of node $i$ (in terms of edges), with $k_{i}=|\partial_{i}|$ and $s_{i}=\sum_{(i,j,w_{ij})\in\partial_{i}}w_{ij}$ the degree and strength of node $i$ respectively. We also let $k_{i}^{(b)}$, $s_{i}^{(b)}$, and $\partial_{i}^{(b)}$ be defined analogously for the backbone network $G^{(b)}$, such that $\partial_{i}^{(b)}=\partial_{i}\cap G^{(b)}$ is the neighborhood of node $i$ in the backbone $G^{(b)}$. Transmitting the network $G$ is equivalent to transmitting $\partial_{i}$ for all nodes $i$, and we can formulate our objective in this way by using the same encoding as in the global case to transmit $\partial_{i}$ using $\partial_{i}^{(b)}$ for all nodes $i$ separately. This simply requires matching up the neighborhood level variables $\{s_{i},k_{i},s_{i}^{(b)},k_{i}^{(b)}\}$ with the global variables $\{W,E,W^{(b)},E^{(b)}\}$ in the description length of Eq. II.1.
Given this correspondence with the global method, the information required to transmit the neighborhood $\partial_{i}$ using the backbone neighborhood $\partial^{(b)}_{i}$ is given by
$\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)})$ $=\log(s_{i}-k_{i}+1)+\log(s_{i}+1)$
$+\log(k_{i}+1)+\log{k_{i}\choose k_{i}^{(b)}}$ (4)
$+\log{s_{i}^{(b)}-1\choose k_{i}^{(b)}-1}+\log{s_{i}-s_{i}^{(b)}-1\choose k_{i}-k_{i}^{(b)}-1}.$
The total description length for the local MDL backboning method is then given by
$\mathcal{L}^{(\text{local})}(G^{(b)})=\sum_{i=1}^{N}\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)}),$ (5)
where $N$ is the total number of nodes in the network.
Since the description length is analogous to the global description length of Eq. II.1, we can use a similar argument as in Sec. II.1 to show that in the regime $s_{i},k_{i}\gg 1$ the neighborhood-level description length $\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)})$ is monotonically decreasing in $s_{i}^{(b)}$ for any fixed value of $k_{i}^{(b)}$, so long as $\frac{s_{i}^{(b)}}{s_{i}}\geq\frac{k^{(b)}_{i}}{k_{i}}$. This implies that a greedy algorithm will also be approximately optimal for constructing the node neighborhoods $\partial_{i}^{(b)}$ within the backbone, and an efficient algorithm for constructing the entire backbone is as follows:
  1. 1.
    Sort the edges $(i,j,w_{ij})\in G$ in decreasing order of weight $w_{ij}$. Initialize the weight-ordered neighborhood $\partial_{i}=\{\}$ for each node $i$.
  2. 2.
    Iterate over the ordered list $G$ and for each edge $(i,j,w_{ij})$, add the edge to $\partial_{i}$. Then, for each node $i=1,...,N$, perform steps 3 - 5 below.
  3. 3.
    Initialize an empty backbone neighborhood $\partial^{(b)}=\{\}$ with $k_{i}^{(b)}=s_{i}^{(b)}=0$ and initialize an array DL to store the description length value for each value of $k_{i}^{(b)}$. Compute $\text{DL}[k_{i}^{(b)}=0]$ using Eq. II.2.
  4. 4.
    Iterate through the ordered edges in $\partial_{i}$, and for each edge $(i,j,w_{ij})$: increment $k_{i}^{(b)}\mathrel{+}=1$ and $s_{i}^{(b)}\mathrel{+}=w_{ij}$ and compute $\text{DL}[k_{i}^{(b)}]$ using Eq. II.2.
  5. 5.
    After adding the last edge from $\partial_{i}$, identify the value $k_{i,\text{local}}^{(b)}=\text{argmin}[\text{DL}]$ for which the description length was minimized, and set the MDL-optimal backbone neighborhood $\partial^{(b)}_{i,\text{local}}$ to be the first $k_{i,\text{local}}^{(b)}$ edges in the ordered $\partial_{i}$.
The above algorithm considers out-neighborhoods $\partial_{i}$, but an analogous procedure holds for in-neighborhoods if we swap $i$ and $j$ in the neighborhood, strength, and degree updates. As with the global method, the bottleneck of this local backboning algorithm is the initial step of ordering the edges by weight, so the method incurs a total runtime complexity of $O(E\log E)$. This is comparable to the complexity of the Disparity Filter serrano2009extracting in practice and more algorithmically efficient than other existing local backboning methods marcaccioli2019polya ; foti2011nonparametric .
We can see that through its information theoretic formulation, the local backboning method applies a weight threshold that adapts to the heterogeneity within each node neighborhood $\partial_{i}$. This will allow it to provide a more compressive backbone for networks with heterogeneous node strengths $s_{i}$ than the global method described in the previous section. To determine which method provides superior compression for a given network $G$, we can run each algorithm and compute the final description length values $\mathcal{L}^{(\text{local})}(G^{(b)}_{\text{local}})$ and $\mathcal{L}^{(\text{global})}(G^{(b)}_{\text{global}})$ using Eq. 5 and Eq. II.1 respectively to identify the lower of the two description lengths.
It is worth noting that one can guarantee that the local MDL backbone constructed using the objective in Eq. 5 is a single (weakly) connected component if all nodes have at least one edge in their neighborhood in the original graph (in the direction considered in the MDL objective). The cases $\{s_{i}^{(b)},k_{i}^{(b)}\}=\{0,0\}$ and $\{s_{i}^{(b)},k_{i}^{(b)}\}=\{s_{i},k_{i}\}$ give equal contributions to the description length, so if the optimal backbone sets $\partial^{(b)}_{i}=\{\}$ we can equivalently set $\partial^{(b)}_{i}=\partial_{i}$ and obtain the same description length. For the experiments in Sec. III we let $\partial^{(b)}_{i}=\{\}$ for these cases, but our code implementation allows for either option. We will see that regardless of this choice the local backbones tend to be more well connected than the global backbones for the networks studied.
It is also useful to construct a normalized inverse compression ratio $\eta\in[0,1]$ to determine the extent to which we can compress a network using its MDL-optimal backbones relative to a naive encoding where we do not use any backbone as an intermediate step (equivalent mathematically to setting $G^{(b)}=\{\}$ as the empty graph). The ratio we use here is given by
$\eta^{(\text{global/local})}=\frac{\mathcal{L}^{(\text{global/local})}(G_{\text{global/local}}^{(b)})}{\max[\mathcal{L}^{(\text{global})}(\{\}),\mathcal{L}^{(\text{local})}(\{\})]},$ (6)
where $G_{\text{global/local}}^{(b)}$ is the MDL-optimal backbone obtained by the global (local) method when computing $\eta^{(\text{global})}$ ($\eta^{(\text{local})}$). We can also choose to normalize Eq. 6 by the description length corresponding to the method of interest (global or local) rather than the maximum, but the max normalization is nice in practice because it allows us to compare (a) the absolute description lengths of the global and local methods and (b) their compression relative to the most naive encoding all in a single measure.
Fig. 1 shows a diagram of both the global and local backboning methods for a small synthetic example network. Code implementing these algorithms can be found at https://github.com/aleckirkley/mdl-network-backbones.
II.3 Bayesian Generalization
It is well known that MDL methods are often equivalent to maximum a posteriori (MAP) estimation in Bayesian generative models mackay2003information . This is also the case for the global and local MDL backboning procedures described in Sec. II.1 and Sec. II.2.
Consider a Bayesian generative model for the weights $\{w_{ij}\}$ of the network $G$ when its edge positions are known. We first assign each edge $(i,j)\in G$ to the (global) backbone $G^{(b)}$ with probability $\pi_{b}$ and to the set of non-backbone edges with probability $(1-\pi_{b})$. This is equivalent to specifying an indicator vector $\bm{b}$ for existence in the backbone such that $b_{ij}=1$ if $(i,j)\in G^{(b)}$ and 00 otherwise. This gives us a prior distribution over the backbone $\bm{b}$ of
$P(\bm{b}|\pi_{b})=\prod_{(i,j)\in G}\pi_{b}^{b_{ij}}(1-\pi_{b})^{1-b_{ij}}.$ (7)
Now assign a weight $w_{ij}$ to the edge $(i,j)$ that is drawn from a geometric distribution with a success parameter $\theta_{b_{ij}}$ that depends on whether we are in the backbone or not (i.e. if $b_{ij}=0$ or $b_{ij}=1$). This allows us to distinguish the weights in the backbone from those outside of it through a mixture model over the edge weights. (There are then two independent parameters, $\theta_{0}$ and $\theta_{1}$.) The likelihood of the weighted network $G$ and the backbone assignment $\bm{b}$ given the model parameters can then be written as
$P(G,\bm{b}|\theta_{1},\theta_{0},\pi_{b})$ $=P(G|\theta_{1},\theta_{0},\bm{b})P(\bm{b}|\pi_{b})$
$=\prod_{(i,j)\in G}[\pi_{b}(1-\theta_{1})^{w_{ij}-1}\theta_{1}]^{b_{ij}}$
×$\times[(1-\pi_{b})(1-\theta_{0})^{w_{ij}-1}\theta_{0}]^{1-b_{ij}}$
@@ II.3 Bayesian Generalization
×$\times\theta_{1}^{E^{(b)}}(1-\theta_{1})^{W^{(b)}-E^{(b)}}$
×$\times\theta_{0}^{E-E^{(b)}}(1-\theta_{0})^{(W-W^{(b)})-(E-E^{(b)})},$
where we have parameterized the geometric distribution so as to enforce $w_{ij}\geq 1$.
Putting a uniform prior $P(\theta_{1},\theta_{0},\pi_{b})=1$ on the model parameters, we can marginalize over these parameters on the unit interval $[0,1]$ to form the marginal joint likelihood
@@
$=B(E^{(b)}+1,E-E^{(b)}+1)$ (9)
$\times$ $B(E^{(b)}+1,W^{(b)}-E^{(b)}+1)$
$\times$ $B(E-E^{(b)}+1,(W-W^{(b)})-(E-E^{(b)})+1),$
@@
Maximum a posteriori (MAP) estimation with this model then aims to solve the following optimization problem to identify the optimal backbone $\bm{\hat{b}}$
$\bm{\hat{b}}$ $=\operatorname*{arg\,max}_{\bm{b}}P(\bm{b}|G)$
$=\operatorname*{arg\,min}_{\bm{b}}\left\{-\log P^{(\text{global})}(G,\bm{b})\right\}.$ (11)
Substituting in Eq. II.3, some algebra shows that
$-\log P^{(\text{global})}(G,\bm{b})$ $=\mathcal{L}^{(\text{global})}(G^{(b)})$ (12)
$~{}~{}~{}~{}~{}~{}+O(\log W+\log E),$
where $\mathcal{L}^{(\text{global})}(G^{(b)})$ is given by Eq. II.1 and the second term contains terms of size less than or equal to $\log W+\log E$. Matching terms with Eq. II.1, we can see that the global description length exceeds $\log W+\log E$ by at least $\log{W-1\choose E-1}$, and so in the limit of $W,E\gg 1$ the fractional error between $-\log P^{(\text{global})}(G,\bm{b})$ and $\mathcal{L}^{(\text{global})}(G^{(b)})$ becomes vanishingly small. In other words, MAP estimation with this Bayesian model is asymptotically equivalent to the MDL minimization problem with the objective in Eq. II.1.
We can generalize this Bayesian model to the neighborhood level for generating the edge weights in $\partial_{i}$, allowing us to make a comparison with the local MDL objective in Eq. 5. In this model, each node $i$ has its own probability $\pi_{b}(i)$ that one of its adjacent edges $(i,j)\in\partial_{i}$ is in the backbone $G^{(b)}$. Similarly, each node $i$ has its own success rates $\theta_{1}(i)$ and $\theta_{0}(i)$ for the geometric distribution over edge weights. We can then compute the joint marginal likelihood $P^{(\text{local})}(G,\bm{b})$ as before by integrating out the parameters $\{\pi_{b}(i),\theta_{1}(i),\theta_{0}(i)\}$ for each node neighborhood $\partial_{i}$ and taking a product over the independent neighborhoods. The result is
$P^{(\text{local})}(G,\bm{b})$ $=\prod_{i=1}^{N}P(\partial_{i},\{b_{ij}\}_{(i,j)\in\partial_{i}})$

### v3

@@ II.1 Canonical Models for Global and Local Backboning
Our problem considers an input network $G=\{(i,j,w_{ij})\}$, which we represent as a set of $E$ weighted tuples $e=(i,j,w_{ij})$. The tuple $e=(i,j,w_{ij})$ indicates an edge from a node $i$ to a node $j$ with weight $w_{ij}\geq 0$ that represents the strength of interaction among $i$ and $j$. The network $G$ must be weighted, as the backboning methods we present find backbones that allow for compression of the edge weights in $G$. The graph $G$ may have self-edges or multi-edges—multi-edges can be combined into a single edge weight—and $G$ may be directed or undirected. In the undirected case we simply consider both $(i,j,w_{ij})$ and $(j,i,w_{ij})$ to be in $G$, with the edge $(i,j,w_{ij})$ appearing in an undirected backbone if either $(i,j,w_{ij})$ or $(j,i,w_{ij})$ is retained in the directed backbone (as done in serrano2009extracting ). Bipartite networks can be accommodated by our framework with no additional modifications. We will let $W=\sum_{e\in G}w_{e}$ denote the total edge weight in the weighted graph $G$. The goal of our network backboning method is to extract a subset $G^{(b)}\subseteq G$ of weighted edges that best compress the overall weight distribution in $G$. We call the subset $G^{(b)}$ the backbone of $G$, which has $E^{(b)}\leq E$ edges with a total weight of $W^{(b)}\leq W$.
To derive a global backboning objective, we can consider a generative process in which we start with an unweighted network $G_{u}$, and for each edge $e\in G_{u}$ we classify $e$ as being either a backbone edge or non-backbone edge by assigning it an indicator variable
$b_{e}=\begin{cases}1,~{}~{}e\text{ is a backbone edge},\\ 0,~{}~{}e\text{ is not a backbone edge}.\end{cases}$ (1)
This results in a binary vector $\bm{b}$ indexed by edge pairs $e=1,...,E$ that partitions the set of edges into a backbone and non-backbone. We then use a mixture model for the edge weights in which the weight $w_{e}$ is generated according to the distribution $P(w_{e}|\bm{\theta}_{b_{e}})$, where $\bm{\theta}_{b_{e}}$ is a set of parameters for the distribution $P$ that depend on whether or not the edge $e$ is in the backbone. We therefore have two independent sets of model parameters $\bm{\theta}_{1}$ and $\bm{\theta}_{0}$ for the backbone and non-backbone edges respectively. This mixture model imposes a distinction between the weight statistics of the backbone and the non-backbone edges, which is the essential feature utilized by the majority of weighted network backboning methods serrano2009extracting ; tumminello2011statistically ; dianati2016unwinding ; casiraghi2017relational ; foti2011nonparametric ; marcaccioli2019polya ; li2007transport .
For the canonical model, we generate the backbone assignments $\bm{b}\sim\text{Bernoulli}(\pi_{b})$, where $\pi_{b}\in[0,1]$ is the probability of an edge being assigned to the backbone. This gives us a prior distribution over the backbone $\bm{b}$ of
$P(\bm{b}|\pi_{b})=\prod_{e\in G_{u}}\pi_{b}^{b_{e}}(1-\pi_{b})^{1-b_{e}}.$ (2)
The likelihood of an observed weighted network $G$ and a backbone assignment $\bm{b}$ given the model parameters and the unweighted network $G_{u}$ can then be written as
$P(G,\bm{b}|$ $\bm{\theta}_{1},\bm{\theta}_{0},\pi_{b})=P(G|\bm{\theta}_{1},\bm{\theta}_{0},\bm{b})P(\bm{b}|\pi_{b})$
$=\prod_{e\in G_{u}}[\pi_{b}P(w_{e}|\bm{\theta}_{1})]^{b_{e}}[(1-\pi_{b})P(w_{e}|\bm{\theta}_{0})]^{1-b_{e}}$
@@
×$\times\prod_{e\in G^{(b)}(\bm{b})}P(w_{e}|\bm{\theta}_{1})\prod_{e\in G\setminus G^{(b)}(\bm{b})}P(w_{e}|\bm{\theta}_{0}),$
where $G^{(b)}(\bm{b})\subseteq G$ is the subset of weighted backbone edges associated with the assignment $\bm{b}$. The above likelihood is (as implied in the model formulation) conditional on knowing the unweighted network $G_{u}$, but we remove explicit dependence on $G_{u}$ in the notation for brevity.
Putting priors $P(\bm{\theta}_{1})$, $P(\bm{\theta}_{0})$, and $P(\pi_{b})$ on the model parameters that are conjugate to the likelihoods $\prod_{e}P(w_{e}|\bm{\theta}_{1})$, $\prod_{e}P(w_{e}|\bm{\theta}_{0})$, and $P(\bm{b}|\pi_{b})$ respectively, we can marginalize over $\{\bm{\theta}_{1},\bm{\theta}_{0},\pi_{b}\}$ to form the joint distribution for the canonical model
@@
$=B(E^{(b)}+1,E-E^{(b)}+1)$
×$\times\int P(\bm{\theta}_{1})\prod_{e\in G^{(b)}}P(w_{e}|\bm{\theta}_{1})d\bm{\theta}_{1}$ (4)
×$\times\int P(\bm{\theta}_{0})\prod_{e\in G\setminus G^{(b)}}P(w_{e}|\bm{\theta}_{0})d\bm{\theta}_{0},$
@@
Maximum a posteriori (MAP) estimation with this model then aims to solve the following optimization problem to identify the optimal backbone assignments $\bm{\hat{b}}$
$\bm{\hat{b}}$ $=\operatorname*{arg\,max}_{\bm{b}\in\{0,1\}^{E}}\left\{P(\bm{b}|G)\right\}.$ (6)
This can be equivalently framed in terms of the description length $\mathcal{L}^{(\text{global})}_{C}(G^{(b)})$ of the canonical global backbone model, which we can parametrize with just the weighted backbone $G^{(b)}$ as
$\mathcal{L}^{(\text{global})}_{C}(G^{(b)})$ $=-\log P^{(\text{global})}_{C}(G,\bm{b})$
$=\log(E+1)+\log{E\choose E^{(b)}}$
$-\log\left[\int P(\bm{\theta}_{1})\prod_{e\in G^{(b)}}P(w_{e}|\bm{\theta}_{1})d\bm{\theta}_{1}\right]$ (7)
$-\log\left[\int P(\bm{\theta}_{0})\prod_{e\in G\setminus G^{(b)}}P(w_{e}|\bm{\theta}_{0})d\bm{\theta}_{0}\right].$
We use the notation $\log(x)\equiv\log_{2}(x)$ for brevity, so that the description length is in units of bits. Eq. II.1 tells us, for an optimal encoding over networks and backbones under our assumptions about the weight distribution, how much information is required to transmit a specific network $G$ and backbone partition $\bm{b}$ mackay2003information .
Solving Eq. 6 is equivalent to minimizing the description length of Eq. II.1, giving the minimum description length (MDL) objective
$\hat{G}^{(b)}=\operatorname*{arg\,min}_{G^{(b)}\subseteq G}\left\{\mathcal{L}_{C}^{(\text{global})}(G^{(b)})\right\},$ (8)
where $\hat{G}^{(b)}$ is the MDL-optimal backbone. We can see that Eq. II.1 has two contributions. The first two terms penalize the complexity of the backbone $G^{(b)}$ by encouraging the backbone size $E^{(b)}$ to be small, for $E^{(b)}\leq E/2$. These terms also encourage the backbone sizes $E^{(b)}$ to be large, for $E^{(b)}>E/2$—the backbone $G^{(b)}$ is equally as compressive as its complement $G\setminus G^{(b)}$ since specifying the backbone edges is equivalent to specifying the non-backbone edges given that the unweighted graph $G_{u}$ is known. (This “bit flip symmetry” under the interchange $0\Leftrightarrow 1$ is intrinsic to information theoretic methods such as mutual information measures felippe2024network .) However, as the backbones $G^{(b)}$ with sizes in the regime $E^{(b)}\geq E/2$ are nearly as dense as the original graphs, they do not provide meaningful sparsification, so we will focus on the regime $E^{(b)}\leq E/2$—it is arguably not meaningful to suggest that more than half of the edges in $G$ have statistically significant weights, regardless of the null model one chooses. The final two terms in Eq. II.1 encourage the weight distributions in the backbone and non-backbone to be distinct and for the model $P$ to provide an effective summary of their variabilities.
We can generalize our canonical Bayesian model to the neighborhood level for the edge weights in each node neighborhood $\partial_{i}=\{(i,j,w_{ij})\}$ for all nodes $j$ towards which $i$ has a directed or undirected edge. We denote the strength $s_{i}$ of each node $i$ as its total incident weight $s_{i}=\sum_{e\in\partial_{i}}w_{e}$ and the degree of node $i$ with $k_{i}=|\partial_{i}|$. In this model, each node $i$ has its own probability $\pi_{b}(i)$ that one of its adjacent edges $e\in\partial_{i}$ is in the backbone $G^{(b)}$. Similarly, each node $i$ has its own model parameters $\bm{\theta}_{1}(i)$ and $\bm{\theta}_{0}(i)$ specific to its neighborhood’s edge weights. This parametrization allows the network neighborhoods $\partial_{i}$ to have different weight distributions, which in turn allows the local backboning procedure to identify statistically significant weights in each neighborhood separately as is done in other local backboning methods serrano2009extracting .
We can then compute the joint distribution $P^{(\text{local})}_{C}(G,\bm{b})$ for this canonical local backboning model by integrating out the parameters $\{\pi_{b}(i),\bm{\theta}_{1}(i),\bm{\theta}_{0}(i)\}$ for each node neighborhood $\partial_{i}$ and taking a product over the independent neighborhoods, assuming a directed network where $(i,j)$ and $(j,i)$ are treated independently. The result is equivalent to assuming $N$ independent copies of the global backboning method, one for each neighborhood $\partial_{i}$, which each have their own associated backbone $\partial_{i}^{(b)}$ with total weight $s_{i}^{(b)}$ and number of edges $|\partial_{i}^{(b)}|=k_{i}^{(b)}$. The resulting joint distribution is given by
$P^{(\text{local})}_{C}(G,\bm{b})$ $=\prod_{i=1}^{N}P_{C}^{(\text{global})}(\partial_{i},\bm{b}_{\partial_{i}})$